OpenStudy (anonymous):

What is the inverse of the function below? f(x) = 11x + 9 f -1(x) = -11x - 9 f -1(x) = x - 9 11 f -1(x) = x - 11 9 f -1(x) = 9 + 11x

OpenStudy (***[isuru]***):

take y = 11x +9 solve it for x 11x = y -9 x = y/11 - 9/11 now switch y an x y = x/11 - 9/11 it's the inverse function u need f^-1(x) = x/11 -9/11 \[f^{(-1)}(x) = \frac{ x-9 }{ 11 }\] got it :) ?

OpenStudy (anonymous):

sorta lol